OpenStudy (anonymous):

What is the radius of a circle in which a 30° arc is 2(pi) inches long? @jh3power would i just multiply 2*pi then work it out like before?

OpenStudy (jh3power):

yeah

OpenStudy (jh3power):

well, but it should be 2pi=30/360 2pi r

OpenStudy (mathstudent55):



OpenStudy (jh3power):

r=radius(answer)

OpenStudy (anonymous):

r being the radius. i would just multiply 2*pi divide 30/360

OpenStudy (jh3power):

so r would be 12

OpenStudy (jh3power):

if it matches the equation i just wrote above, you're right.

OpenStudy (anonymous):

this confuses me way to much

OpenStudy (jh3power):

so...theta/360 times 2 pi r = arc

OpenStudy (jh3power):

r=arc times 360/theta time 1/2pi

OpenStudy (anonymous):

well 2pi is 6.28, and 30/360 is like .0833. multiply those two i get .5

OpenStudy (mathstudent55):

The length of an arc, s, is: \(s = \dfrac{\theta}{360^o} 2\pi r\) where \(\theta\) = degree measure of arc. Here, \(\theta = 30\) and \(s = 2 \pi \), so \(2 \pi = \dfrac{30^o}{360^o} 2\pi r\) \(1 = \dfrac{1}{12} r\) \(r = 12 \)

OpenStudy (jh3power):

you should use my r=~ equation.

OpenStudy (jh3power):

do you get it now?

OpenStudy (jh3power):

don't be confused with r and arc..

OpenStudy (anonymous):

how do you get 1, when 2pi is 6.28? and yeah i do thanks guys

OpenStudy (jh3power):

again, look at my equation. 2pi times 1/2pi times 360/30

OpenStudy (jh3power):

equals r

OpenStudy (jh3power):

because arc=2pi

OpenStudy (anonymous):

is that .5pi times 360/30?

OpenStudy (jh3power):

no... 2pi times 1/2pi =1 !!

OpenStudy (jh3power):

and it makes r=360/30

OpenStudy (jh3power):

and to remind you...don't substitute for pi until you get the whole equation. it only makes it more confusing.

OpenStudy (anonymous):

okay. i get it now,

OpenStudy (jh3power):

welcome